The distance traveled at a constant speed is directly proportional to the time of travel. If Olivia traveled 112 miles in 3.5 hours, how many miles will Olivia travel in 8.9 hours at the same constant speed?
Answer by stanbon(75887)

The distance traveled at a constant speed is directly proportional to the time of travel. If Olivia traveled 112 miles in 3.5 hours, how many miles will Olivia travel in 8.9 hours at the same constant speed?
-----------------
Proportion:
x/8.9 = 112/3.5
---
x = 8.9(112/3.5)
x = 284.8 miles
=====================
Cheers,
Stan H.